### v2

# Optimal market-neutral currency trading on the cryptocurrency platform

Abstract
This research proposes a novel arbitrage approach with respect to multivariate pair trading called Optimal Trading Technique (OTT). We introduce the method to selectively form a ‘bucket’ of fiat currencies anchored to cryptocurrency for simultaneously monitoring and exploiting trading opportunities. To handle the quantitative conflicts that arise when receiving multiple trading signals, a bi-objective convex optimization process is designed to cater to the investor’s preference between profitability and risk tolerance. This process includes tunable parameters such as volatility punishment, action thresholds. During our experiments in the cryptocurrency market from 2020 to 2022 when the market was experiencing a vigorous bull-run immediately followed by a bear-run, the OTT realized an annualized profit of 15.49%. We further carried out the experiments in bull, bear, and full-cycle market conditions separately, and found that OTT is capable of achieving stable profit under various market conditions. Apart from the profitability side of the OTT, the arbitrage operation provides a new perspective of trading, which requires no external shorting and never hold intermediate cryptocurrency during the arbitrage period.
keywords:
Pair Trading; Multivariate; Quantitative Trading; Cryptocurrency Market
articletype: RESEARCH ARTICLE
1 Introduction
The secondary market is a modern, efficient form of double auction. Participants with the intention to trade assets bring their acceptable price to bid and ask. The highest bid and lowest ask are to be matched if the price meets. At the macro level, we believe in the efficiency of the secondary market (Fama 1970). Nevertheless, countless economic bubbles bursting in history tells us that investors do not always act rationally. Dale et al. (2005) in the story of South Sea Company depicted a picture of how an orderbook filled with irrational bids and asks could push the price in the wrong direction. Finding out the inefficiencies in an efficient market maintains the equilibrium of market trading. Market directional trading is astute traders estimate a ‘more correct’ pricing for traded instruments and give orders with expectation for the market to revert to being rational. Long stands for traders spend their money to buy the asset for betting on the market to go up and sell later; Short on the opposite means traders borrow the asset and sell to the market to hold money until the market stops dropping to buy back at a cheaper price. Traders that bet in the right direction are rewarded by the market mechanism for correcting the errors. It is a part of the price discovery process. Instead of scrutinizing the market with the help of experienced accountants or financial analysts, we utilize statistics and optimization to dive into applying quantitative trading.
The recent trending cryptocurrency market benefits from the latest technology advancement and global participation. Its volatility, liquidity and 24/7 trading, make the cryptocurrency market an ideal marketplace to conduct algorithmic trading. Since a user can freely trade between cryptocurrencies and fiat currencies within this market, the cryptocurrency market’s volatility comprises the volatility of forex and cryptocurrencies. If we take advantage of the inefficiencies in the emerging cryptocurrency market and capture anomalous movements from the movements of cryptocurrencies, we shall be able to generate profit from the fast-paced market.
The method of trading we focus on in this research is Pair trading. It is one of the most widespread strategies based on the ideology — ‘Law of one price (LOP)’, which suggests that homogeneous items should ideally have the same price everywhere. However, it is often violated because of real world restrictions such as trade frictions and legal regulations (Isard 1977; Miljkovic 1999). That means, though similar assets in a pair should have similar price trends, their pricing independence can sometimes lead to divergence. Betting on the direction of convergence often produces profit. Trading approaches based on LOP are called Arbitrage and are perceived to have more stable profitability. The stability of profiting is favored by investors. Therefore there have been decades of history for investment banks and hedge funds to utilize pair trading as part of their strategies (Gatev et al. 2006; Elliott et al. 2005; Burgess 2000). It is common to find the trading strategy, as the name describes, in pairs. While there has been a history of applying pair trading techniques in real life, we do not yet have many applications for expanding pair trading to multivariate trading. This paper shall take a deep dive into the pair trading technique and introduce an optimization-based technique to ‘pair trade’ multivariate assets and experiment in cryptocurrency market.
This research has multiple novel contributions: ① Unlike classical techniques (Perlin 2007; Galenko et al. 2012; Dunis and Ho 2005), the proposed approach uses optimization to arbitrage multiple pairs simultaneously to the individual asset granularity without grouping assets. ② The presented bi-objective optimization is adjustable to cater to investors’ preferences by the inclusion of tuneable parameters such as risk tolerance, operating thresholds for maximizing return and minimizing the risk. ③ Our arbitrage approach never holds the volatile intermediate cryptocurrency throughout the arbitrage period. By avoiding the hold of cryptocurrency, the position of the trading system consists of only fiat currencies, hence lower risk. ④ The proposed trading approach is designed in a way that no borrowing is required during shorting. All the trading actions are achievable within the basket of the initial principal. Therefore, we save on the cost and hassle of external borrowing. ⑤ Lastly, the proposed technique displays a strong robustness of market-neutrality and profitability in both bull and bear markets.
The remainder of the article is structured as follows. Sections 2 review the most cited works in pair trading and give the crux of the proposed trading idea. Sections 3 introduce the pre-requisite knowledge and the details of the proposed trading methodology, respectively. Simulation algorithm is presented in Section 4. Finally, results and conclusions are presented in Sections 5 and 6, respectively.
2 Related Work
2.1 Pair Trading
Long-Term Capital Management L.P. (LTCM)was a highly-leveraged hedge fund founded by John Meriwether in 1994 operated on bond pair trading. It is famous for its billions of dollars in bailout after the 1997 financial crisis and is known as an early real-life example of pair trading (Slivinski 2009).
Figure 1: Pair trading between assets y1 and y2
The detailed operation of pair trading is rather simple. Figure 1 is a pedagogical demonstration based on the most cited baseline pair trading framework (Gatev et al. 2006), which is also known as Distance Method. The Distance Method of pair trading can be summarized into three steps.
  1. 1.
    Exhaustively search for two assets that have the closest price movement in history as y1 and y2. Group y1 and y2 as a trading Pair.
  2. 2.
    Difference the price of y1 and y2 as Distance. Set position opening threshold as two times of Distance’s historical standard deviation (presented as the band area with y1). Once the price of y2 surpasses the threshold, open a position by shorting y2 and longing y1. Abide time until the price of y1 and y2 crosses, then close the position by reversing the buy and sell actions from the position opening.
  3. 3.
    If y2 goes below the threshold, do the opposite by longing y2 and shorting y1 until the next crossing point. Repeat the practice till end of the trading period.
A critical property of pair trading is that we can achieve Market Neutrality if we carefully maneuver our long and short positions. Market neutrality is the property that the trades are profitable irrespective of the direction the complete market, or the underlying assets being traded move towards. Market neutrality is obtained by a simple principle: if the distance between two correlated assets diverges in the trading period, bet the equivalent quantity on the long side and the short side. The criteria of the divergence can also be versatile. In the work of Chen et al. (2018), the authors form pairs from asset returns instead of their prices, based on the idea that a temporarily weaker return implies a stronger rebounce (and vice versa). From another perspective, both Gatev et al. (2006) and Chen et al. (2018) construct their work based on correlation.
Besides correlation, there is an essential concept in the time-series analysis called cointegration. Cointegration is defined by the co-movement of two time-series data sets and is believed to represent a long-run relationship (Engle and Granger 1991). In the book by Vidyamurthy (2004), he examined assets for cointegration with the Engle-Granger test and bound assets based on not only correlation but also cointegration. His work provided good guidance for cointegration-based pair trading.
A step ahead, there are many studies discussing multivariate pair trading. They adopt the idea of hedging market risk by managing positions and creating ‘artificial pairs’ from scenarios with more than two assets. The technique developed by Perlin (2007) combined assets into an artificial index to form a pair. The main idea is that, instead of finding a pair, he created a synthetic index $P_{it}^{*}$ to anchor a real index $P_{it}$. The synthetic index $P_{it}^{*}$ could be constructed by giving weights ($w$) to $m$ number of assets based on their prices:
$P^{*}_{it}=\sum_{k=1}^{m}w_{k}P_{kt}^{*}$
We can tune the weights among the participating assets so as to match them close to our asset $P_{it}$ as much as possible. The next step is to arbitrage $P_{it}$ and $P_{it}^{*}$ for pair trading. The methodology of artificial pair exists in cointegration methods as well. Galenko et al. (2012) used a moving window method to pair trade among a group of assets. In the moving window, the authors calculated the cointegration vector $\boldsymbol{b}$ to track $Z_{t}$ from the asset return combined with cointegration value $b$.
$Z_{t}=\sum_{i=1}^{N}b^{i}r_{t}^{i}$
Next, assets were partitioned into two categories by the results (positive or negative) of $\boldsymbol{b}$. In the two categories: $L$ (Long) and $S$ (Short), quantified $Z_{t}$ could provide quantifiable recommendations for how much should be invested in each asset. A similar methodology is adopted by Dunis and Ho (2005). They track the EURO STOXX50, and split the outperformers and underperformers into STOXX50+ and STOXX50-. Next, they pair traded STOXX50+ and STOXX50- to achieve ‘double alpha’. Because this method took on non-negligible market risk, therefore it is a weak form of pair trading. We can see that the majority of the existing pair trading techniques are centered around the idea of compounding pairs. In our research, we aim to provide a technique that neither makes an artificial index nor binds a group of assets. With the inclusion of optimization, we can operate directly on all participating assets.
2.2 Cryptocurrency
The work in Fang et al. (2022) provides a comprehensive survey of the history and recent research about cryptocurrencies. Cryptocurrency was started by the groundbreaking work of Nakamoto (2008) to introduce a decentralized monetary system. The property of decentralization comes from the idea of a peer-to-peer ledger recording system. By incentivizing Bitcoin users to record a full transaction ledger on the user’s own computer, the ledgers are distributed in countless individual nodes instead of a centralized server. The property brings frictionless international participation of Bitcoin. Furthermore, global participation provides excellent liquidity for Bitcoin trades. Not soon after, the success of Bitcoin was followed by imitators. Ethereum (ETH) is the second most famous cryptocurrency in the queue created by Buterin (2014). Ethereum is claimed to be the next generation of blockchain because of its programmable extensibility. In the application, ETH is our quote currency for arbitrage. We choose ETH because of the accessibility of data and high liquidity.
Popular pair trading techniques in Section 2.1 are also introduced in the Cryptocurrency market. The work in Fil and Kristoufek (2020) had a distance method application based on the work of Gatev et al. (2006), which achieved 3% monthly profit in the Cryptocurrency market. The work in Van den Broek and Sharif (2018) formed pairs based on 60-day trading periods to profit on arbitrage opportunities and proved the inefficiency of the cryptocurrency market.
Figure 2: The price of ETH price in multiple currencies
Similar to gold, ETH has a global market in various currencies. An ETH can be traded independently with most major fiat currencies. Trading behavior incurs volatilities, and the volatilities of ETH in different fiat currencies are not always synchronized. The tiny gaps of inefficiencies are the trading opportunities. What makes ETH a better anchor than gold is that ETH as a cryptocurrency is completely fungible and has low international transaction frictions. As we can see in Figure 2. Though the CAD, EUR, GBP and USD follow a similar trend, they do not have a 100% correlation. Our technique expects to capture the minor volatility movement and arbitrage the market inefficiencies.
Though the cryptocurrency system functions in a decentralized manner, cryptocurrency users can deposit their Bitcoin or Ether into a centralized exchange for trading. The exchange works just the same way as stocks or forex exchange. Where you can freely exchange your cryptocurrency for another cryptocurrency or fiat currency, on spot or margin trading. Due to the decentralized nature of participants on the cryptocurrency exchange, it functions 24/7 with liquidity coming from global participation. Due to these properties, there have been experiments on the cryptocurrency exchanges since as early as 2014 (Shah and Zhang 2014; Brandvold et al. 2015; Cheah and Fry 2015).
3 Methodology
There are two stages in our trading technique. The first stage is a screening of the assets. We select the assets that are correlated and cointegrated with each other to participate in the asset trading bucket. Once the bucket is formed, we monitor the difference between each asset and conduct trades for opening/closing trading signals.
3.1 Screening
Figure 3: The volatility of forex rate against USD
In the screening stage, we need to analyze the historical data so that the assets are suitable for arbitrage. Since the movement of cryptocurrency is impacted by the volatility of both ETH and fiat currencies, if a fiat currency suffers from severe inflation or deflation, the value change of that fiat could mislead our trading signals. Hence, one of the major participating criteria is the volatility of fiats should be small compared to the cryptocurrency. Figure 3 is an illustration of strong fiat depreciation. The difference between AUD and NZD is relatively stable compared to the rapid drop of JPY. If our technique is trading within the period from Mar 2021 to Sept 2022, the depreciating JPY will send a short signal out of its value change with no mean reversion. To select appropriate fiat currencies into the trading bucket, we reviewed the historical trend of fiat currencies to make sure that they have desired statistical relationships.
The first desired statistical property is correlation. We examine the correlation by Pearson’s correlation coefficient (Cohen et al. 2009). Pearson’s correlation coefficient ranges between $[-1,1]$. The higher the absolute value of the coefficient, the stronger association between the two examees. -1 and 1 represent an absolute negative and positive association, and 0 in the between is for no association between datasets.
The second desired statistical property is cointegration. We call the price difference between two assets as spread. In the long term, a widening spread exposes arbitrageurs’ position to bankruptcy risk. That is the incentive to expect the spread between every two assets to have a constant mean and variance. The characteristic of a time-series dataset with a constant mean and variance is called stationary. Cointegration, is the relationship to call the two time-series datasets whose difference is stationary. Cointegration is vital in pair trading because when we form a pair with cointegration in the cryptocurrency market, longing one asset and shorting another is effectively equivalent to holding the price difference. If the spread is stationary with a constant mean, a convergence should always follow a divergence. And the property is called mean-reversion. In this research, the Eagle-Granger test (Engle and Granger 1987) is used to examine the cointegration relationship.
3.2 Long and Short Trading Signals
After we have selected the assets, we need to set up rules for how and when to trade the multivariate assets. Some questions ahead of us are: How do we ensure market neutrality for multiple assets? How should we define and react to the trading signals? How do we allocate our principal capital?
3.2.1 Multivariate Pairs
We want to give a definition and a consistent notation for Pairs. The assets in this research are fiat currencies against ETH such as ETH/USD, ETH/CAD, ETH/GBP. When a pair between ETH/USD and ETH/CAD is formed, the pair will be called as ETH/USD:ETH/CAD, which can be abbreviated to USD:CAD. Theoretically, we can maximum form $k(k-1)/2$ pairs from $k$ assets.
3.2.2 Buy/Sell Signaling
Arbitrage relies on anomalies. We extract anomalies by comparing normalized spread with standard deviation. We have Open_Threshold and Close_Threshold, the two selected parameters for tuning the trading signals. For a two-asset trading scenario: The trading process between the assets $i$ and $j$ can be generated as follow:
  • Step 1: Logarithmize the Price of $i$ and $j$ to reflect the exponential price growth in the cryptocurrency market.
  • Step 2: Difference logarithmized Price as the Spread array.
  • Step 3: Extract the Standard Deviation from Spread, and normalize the spread array by z-score.
  • Open: If there is not a position opened already, and z-score $>$ Open_Threshold $\times$ Standard Deviation, then open a new position.
  • Close: If there is an open position, and z-score $<$ Close_Threshold $\times$ Standard deviation, then close the position.
  • Step 4: Repeat Open and Close steps until the end of arbitrage.
We give a graphical example to explain the workings of the proposed approach. Table 1 gives the statistics and graphs for a three-asset scenario. Since each asset moves independently, we can see that each pair has a different standard deviation after normalization. The spreads are volatile, and we draw thresholds in auxiliary lines. Every touch of the Open_Threshold represents a position opening signal. On the contrary, if an open position exists between the spread and the spread touches the Close_Threshold, it sends out a position closing signal.
Table 1: Spreads trading signaling
Spread Standard Deviation Open_Threshold Close_Threshold
USD:CAD 0.024017 0.168122 0.144102
USD:GBP 0.059545 0.416817 0.357270
CAD:GBP 0.049811 0.348677 0.298866
  • In a cut-off period on 17 Jan 2021, a visualized spread trading signalling based on the OPEN_THRESHOLD as 7 and CLOSE_THRESHOLD as 6.
3.2.3 Position Movements
Suppose we get back from the spreads to the actual asset granularity (Figure 4). The initial principal is distributed amongst the fiat currencies we are trading. Traditionally, we long the loser and short the winner for every positive signal. However, under multiple spread conditions, the trading signals are calculated collectively at every trading interval. The subplot 4(a) displays the ETH price after normalization as price index to express a comparable price growth of ETH in USD, CAD and GBP, and the subplot 4(b) tracks the principal movement of our fiat currencies.
Figure 4: Actions and position movements for a three-currency scenario
We can see that at time $t_{1}$, GBP starts a rapid decline. Along with that, both relevant spreads GBP:CAD and GBP:USD are shrinking. This is the time ($t_{1}$) for trading signals from spread GBP:CAD and GBP:USD. Our corresponding actions are longing ETH with GBP, and shorting the ETH to USD and CAD. At the latter time $t_{2}$, there is a rebound in ETH price against all currencies. Though all fiat currencies are going up, the USD reversal is weaker than the other two, which makes USD:CAD and USD:GBP spread greater than the open threshold. It triggers the USD to be spent for buying ETH and selling the ETH for CAD and GBP.
By design, it is possible that a fiat currency needs to be longed and shorted at the same time. Besides, the classical scenario of two-asset trading still exists because each spread is monitored individually. An example if the movement that happened at $t_{3}$ does not trigger trading in CAD. The latency of rebound in USD makes only the $z$-score of USD:GBP greater than open thresholds. Therefore the position only opened for the spread USD:GBP. Later after $t_{4}$ when the spread widens, the position opened at $t_{3}$ is closed.
There are more benefits of multivariate asset inclusion. On the one hand, if a fiat currency is receiving a long signal from one pair and the same amount of another short signal, the action can be canceled off on paper to save fees. On the other hand, when one fiat currency reveals a strong anomalous movement, all relevant pairs shall display signals for the principal flowing in/out from the fiat currency. If there is an entry signal in USD:GBP, all the USD can be used to conduct the operation. There will be no USD sitting on the bench and missing out. Though we like to see the perfect mitigation saves on the transaction fee. Signals that can cancel off on each other are rare cases. More often, we have multiple short and long signals that happen simultaneously. And it may receive one short signal from a position closure and another short signal from a new position opening. Therefore, we need to devise a mechanism to determine how exactly the funds should be distributed.
3.3 An Optimal Approach to Capital Allocation
Resource distribution is a question about how much to invest in each asset every time when there is a trading opportunity. Therefore, at every trading interval, as long as a trading opportunity is detected, the optimization process is executed for a quantitative decision on the amount to invest. We don’t need optimization if we only ever trade a two-asset pair since all the capital (or a pre-determined amount) can be allocated for the long and short positions to be taken. When we have more than two assets, the multilateral relationships require that we ensure the resources distributed on each signal result in the best outcome. The optimization process aims to maximize return with minimum risk under restrictions. In this research, a bi-objective optimization formulation inspired from the Markowitz portfolio allocation (Markowitz 1952) operates to find the optimal investment solution.
3.3.1 Variable
Let Weights $w\in\mathbb{R}^{N\times 2}$ be the variable of the optimization in the form of a matrix for weights from every pair. The first column is the long currencies with positive weights, and the second column is the short currencies with negative weights. The negative weights capture the fact that the capital reduces when going long, and the capital increases when going short. An example of a weight matrix is shown below:
$\left[\begin{array}[]{c c}w_{USD}&w_{CAD}\\ w_{USD}&w_{GBP}\\ w_{CAD}&w_{GBP}\\ \end{array}\right]$
3.3.2 Parameters
  1. 1.
    Let $\mathbf{C}$ be a vector of participating currencies, where $|\mathbf{C}|\in\mathbb{Z}^{+}$ represents the number of currencies.
  2. 2.
    Let $\mathbf{N}$ the a vector of Pairs formed from Currencies, where $|\mathbf{N}|\in\mathbb{Z}^{+}$ represents the number of pairs.
  3. 3.
    Let $\mathbf{t}\in[1,2,3,...,T]$ represent the timestamp of execution, and $\mathbf{T}$ denotes the total length of the arbitrage period.
  4. 4.
    Let $\mathbf{P}\in\mathbb{R}^{|C|\times T}$ be a matrix representing ETH prices in $|C|$ fiat currencies for the arbitrage period $T$.
    $\begin{cases}\text{ETH/USD}:[P_{USD,t=1},P_{USD,t=2}...,P_{USD,t=T}]\\ \text{ETH/CAD}:[P_{CAD,t=1},P_{CAD,t=2}...,P_{CAD,t=T}]\\ \text{ETH/GBP}:[P_{USD,t=1},P_{USD,t=2}...,P_{USD,t=T}]\\ \end{cases}$
  5. 5.
    Let $\mathbf{r}\in\mathbb{R}^{|C|}$ recording the average returns of ETH price in $|C|$ fiat currencies throughout the arbitrage period and let Expected Profit $\mathbf{EP}\in\mathbb{R}^{|N|\times 2}$ be a matrix for $N$ pairs recording how much profit is expected on average once an arbitrage position is opened. Because each spread consists of two different assets, the speeds of mean-reversion should differ from pair to pair. Hence, we measure a parameter vector $\mathbf{mr}$ of type $\mathbb{R}^{|N|}$ to record the average reversion time to the mean from historical data. Expected Profit is calculated based on the combination of Returns and Mean-reversion speed as follows:
    $\left[\begin{array}[]{c c}mr_{USDCAD}\times r_{USD}&mr_{USDCAD}\times r_{CAD}\\ mr_{USDGBP}\times r_{USD}&mr_{USDGBP}\times r_{GBP}\\ mr_{CADGBP}\times r_{CAD}&mr_{CADGBP}\times r_{GBP}\\ \end{array}\right]$
  6. 6.
    Let Covariance $\mathbf{cov}\in\mathbb{R}^{N\times 2\times 2}$ be a three dimensional matrix. It contains $2\times 2$ matrices for $N$ pairs. The covariance matrix is used to compute the risk of opening this position. An individual $2\times 2$ matrix from the three-dimensional slice is a variance-covariance matrix for the two fiat currencies in a pair. An example of the $2\times 2$ looks like this below:
    $\begin{bmatrix}\text{Var}(USD)&\text{Cov}(USD,CAD)\\ \text{Cov}(USD,CAD)&\text{Var}(CAD)\end{bmatrix}$
  7. 7.
    Let Transaction Cost $\mathbf{tc}\in\mathbb{R}^{\geq 0}$ be a constant, which represents the percentage to be deducted from each investment amount as commission.
  8. 8.
    Let Trading Weights $\mathbf{TW}\in\mathbb{R}^{|C|}$ be a vector of currency weight occupation for every currency. Since our arbitrage is a continuous process, every trade decision carries historical information. A strong buy signal is guiding the optimization process to invest all the USD into ETH. Because we do not use margin trading, hence the maximum we can invest is 100%. However, if 40% USD has been used to long USD:CAD in an unclosed open position, the maximum we can invest for the next signal falls down to 60%. In this example, the 40% is carried from a weight that is being traded. That is why we give it the name: Trading Weights (TW), and it can be positive or negative.
3.3.3 Objective
The objective of the optimization is to maximize the return while minimizing the risk. We receive trading signals from spreads, so the decisions for investment weights are in pairs. When we need to long USD and short CAD in the pair USD:CAD, we are effectively applying the weight to action on betting the rise of $r_{USD}$ and the drop of $r_{CAD}$. The expected profit is weights multiply the expected profit presented as:
$\sum_{i=1}^{|N|}w_{n}\cdot(EP_{n}\odot[1,-1])^{{}^{\prime}}\,,\,\forall n\in N$ (1)
Where $\odot$ is the Hadarmard product, as the long weights are positive and short weights are negative. We also expect to minimize risk, which means the portfolio has to have as little volatility as possible. According to the modern portfolio theor (Markowitz 1952), the portfolio volatility is $\boldsymbol{w}\times\boldsymbol{cov_{2\times 2}}\times\boldsymbol{w}^{{}^{\prime}}$ for any pair spread in $N$. Once again, as the long weights are positive and short weights are negative, we need to take the Hadamard product with $\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}$ to correctly capture the portfolio covariance in the covariance-variance matrix. Overall, for $N$ pairs we get Equation (2).
$\sum_{i=1}^{|N|}w_{n}\cdot cov_{n}\odot\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}\cdot w_{n}^{{}^{\prime}}\,,\,\forall n\in N$ (2)
3.3.4 Weight Constraints
The first constraint is that we can never invest more than what we have. Therefore, in every trading pair, the long weights cannot surpass 100%, and the short weights shall not be less than -100%:
$0\leq w_{n,long}\leq 1,\;-1\leq w_{n,short}\leq 0,\;\forall n\in N\$ (3)
Another constraint comes from the weight at the currency level. Every currency should not be spent more than 100% at any time of the trade. Every time the optimization process runs, the process calculates the summation of current trading weight, long weight and short weight to ensure that it does not exceed 100% for each individual currency.
$TW_{c}+\sum_{n=1}^{N}w_{c,long}-\sum_{n=1}^{N}w_{c,short}\leq 1\,,\,\forall c\in C$ (4)
3.3.5 Market Neutrality
Market Neutrality is the ability to generate stable income regardless of the market condition, and it is one of the major incentives for investors to select pair trading from other quantitative trading methods. To achieve neutrality, we have to ensure that the equivalent amount of ETH in long and short positions are always equal. In that, we need to translate the weights back to the amount of ETH. The price of currency $c$ at time $t$ is attainable as $P_{c,t}$.
In our case, whenever an ETH is bought in one currency, that ETH should be sold instantaneously. Hence, we expect the last constraint is that every pair position reflecting back on assets is equalized in the unit of ETH. Transaction cost tweaks the balance slightly. The amount of ETH received from longing is reduced because of transaction cost. However, the amount of ETH required to get a certain amount of fiat currency increases. The market neutrality constraint is presented in Equation (3.3.5):
$-w_{n,short}\cdot(1+tc)P_{c1,t}=w_{n,long}\cdot(1-tc)P_{c2,t}$
$\forall t\in T,\forall(c1,c2)\in C\;,\forall n\in N$ (5)
3.3.6 Optimization Equation
Given the above, for the purpose of maximizing the profitability and minimizing the risk, one can optimize the allocation of the principal capital by solving the bi-objective optimization problem in Equation (6). In the optimization process, the ultimate goal is to calculate a set of quantitative weights when opening any position on the pair spreads.
We understand that market participants have different tendencies toward risk aversion. Therefore, we use a coefficient $\lambda$ (default to 1) to measure the degree of risk aversion in the objective. The greater the $\lambda$, the greater the risk aversion. In extreme cases, $\lambda=0$ stands for risk neutrality. The $\lambda$ and portfolio variance form a punishment factor in the objective to cater for investors’ risk preferences.
$\begin{cases}\text{obj}:&\sum_{i=1}^{|N|}w_{n}\cdot(EP_{n}\odot[1,-1])^{{}^{\prime}}\,,\,\forall n\in N\\ &-\lambda\sum_{i=1}^{|N|}w_{n}\cdot cov_{n}\odot\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}\cdot w_{n}^{{}^{\prime}}\,,\,\forall n\in N\\ s.t.&0\leq w_{n,long}\leq 1,\;-1\leq w_{n,short}\leq 0,\;\forall n\in N\\ \\ s.t.&TW_{c}+\sum_{n=1}^{N}w_{c,long}-\sum_{n=1}^{N}w_{c,short}\leq 1\,,\,\forall c\in C\\ s.t.&-w_{n,short}\cdot(1+tc)P_{c1,t}=w_{n,long}\cdot(1-tc)P_{c2,t}\\ &\forall t\in T,\forall(c1,c2)\in C\;,\forall n\in N\end{cases}$ (6)
4 Simulation
This section gives the process and structure of our simulation (Algorithm 1) on historical data. The process contains multiple stages — screening, hist_analysis, trading and optimisation.
Data: Pricing_Data
/* Historical Pricing data for ETH against USD, EUR, CAD, GBP */
1 def optimisation(trading_data, expected_return, expected_risk ):
2       invest_weights $\leftarrow$ trading_data, expected_return, expected_risk
       /* Optimize investment weights from historical data using Equation (6) */
3       return invest_weights
4
5def trading(open_date, close_date, trading_data, open_threshold, close_threshold, expected_return, expected_risk) :
6       orig_amount()
       /* Set the starting amount of each asset */
7       spread_data $\leftarrow$ trading_data
       /* calculate spread data from assets trading period pricing data */
8       for time in [open_data, close_date] do
[…]
10                   close_position()
                   /* Close position for a pair when there is a position opening and the spread is smaller than the closing threshold */
11
[…]
13                   invest_weights $\leftarrow$ optimisation(trading_data, expected_return, expected_risk)
14                   open_position(invest_weights, orig_amount)
                   /* Open position when the spread has no current position and is greater than the opening threshold */
15
16
17
18def hist_analysis(formation_data, trading_data ):
19       std_dev, expected_return, expected_risk $\leftarrow$ formation_data
       /* Calculate standard deviation, expected risk/return among spread */
20       open_threshold, close_threshold $\leftarrow$ std_dev, formation_data
       /* Set trading threshold between spreads */
21       trading(open_date, close_date, trading_data, open_threshold, close_threshold, expected_return, expected_risk)
       /* Call trading function to trade from open_date to close_date */
22
23def screening(pricing_data):
24       formation_data, trading_data $\leftarrow$ pricing_data
       /* Split the data into formation period and trading period */
25       assets_correlation, assets_cointegration $\leftarrow$ formation_data
       /* calculate the correlation and cointegration between assets */
26       participating_assets $\leftarrow$ assets_correlation, assets_cointegration
       /* Select assets above threshold */
27       formation_data, trading_data $\leftarrow$ formation_data, trading_data, participating_assets
       /* filter the data to only participating assets */
28       hist_analysis(formation_data, trading_data)
29
screening(Pricing_Data)
Algorithm 1 Arbitrage Technique Simulation
The screening (c.f. Section 3.1) determines what assets should be included within our trading bucket. ETH pricing data is extracted from Crypto exchanges against fiat currencies as the input (Algorithm 1 Line 1). We choose which period we trade upon by splitting the pricing data into formation data and trading data. In the formation period, ETH prices against fiat currencies are screened for filling the bucket of fiat currencies with strong correlation and high cointegration test passing rate. Once we dropped the fiat currencies with weak correlation/cointegration relationships, the screened currencies are included in $participating\_assets$. Then we call function hist_analysis.
In the hist_analysis process (Algorithm 1 Line 1), standard deviation, expected return, and expected risk are computed using the historical formation data. Parameters OPEN_THRESHOLD and CLOSE_THRESHOLD are computed using the formation data (c.d. Section 3.2.2). Knowing when to trade is not sufficient enough, we also need to calculate how much to bid/ask on each order. The trading function (Algorithm 1 Line 1) simulates by amount. The first step is to set up $orig\_amount$, which is how much for participating fiat currency we have at the beginning of the trade. Secondly, every two currencies are combined into spreads. Next, we loop through all the trading times and constantly check if any open position meets the closing criteria:
  • if yes, we close the open position to harvest profit, if any. Trading weight vector $TW$ is also updated as described in Section 3.3.
  • if no, we leave it as is and check again in the next time sample
Same as the closing, all the spreads that do not have an ongoing open position need to be checked if an open signal hits at every time sample. Different from closing a position, open-position actions must be considered for how much $orig\_amount$ to spend in each open signal.
To calculate the amount to invest, we call the optimisation function (Algorithm 1 Line 1) by feeding trading data and properties of trading data such as expected return and expected risk to calculate the weights. The optimization is based on Equation 6. Once the optimal weights are calculated, it is returned to the trading function for the continuation of position opening till the next call.
5 Experimental
In this section, we describe the experiments carried out on historical datasets. We first describe the experimental dataset and setup followed by the result of applying the proposed Optimal Trading Technique (OTT)(c.f. Algorithm 1) in different market conditions.
5.1 Experimental Setup
We trade a bucket of fiat currency pairs referenced to ETH for a number of reasons: ① Fiat currency pairs (equivalently spreads) referenced to ETH are cointegrated and stationary; ② Cryptocurrency markets have substantial liquidity and depth; ③ The volatility in the cryptocurrency market overwhelms the volatility inherent volatility in the fiat, which gives us opportunities to arbitrage. We adopt the trading bucket in the ETH pairs on the Kraken [^1] https://www.kraken.com/ exchange, where the assets are freely accessible to all participants and yet traded independently. We select the cryptocurrency exchange Kraken because it is one of the top exchanges providing most fiat options for platform users without geographical restrictions. Following this practice, we are adopting mainstream fiat currencies with considerable market activity toward cryptocurrency and have been traded for a substantial period of time.
The datasets we use are publicly available from Kraken OHLCVT dataset [^2] The dataset is downloadable at: https://support.kraken.com/hc/en-us/articles/360047124832-Downloadable-historical-OHLCVT-Open-High-Low-Close-Volume-Trades-data, which is the historical trading data from Kraken exchange for Open price, Highest price, Lowest price, Close price, trading Volume and number of Trades at different sampling time intervals. The sampling time intervals adopted in this experiment are 1 min, 5 min and 60 min.
5.1.1 Accounting for transaction fees
Transaction cost in quantitative trading is always a non-negligible factor to be considered. One of the incentives to introduce optimal trading is the benefit of reducing transaction costs. Buy-and-hold strategy receives the least impact from transaction costs.
Kraken exchange sets a volume-based fee scheme. The more trading volume, the lower the transaction fee can be. In practice, quantitative trading normally involves large trading volumes. We adopt a flat 0.1% transaction fee for long and short orders.
5.2 Data Source Screening
As mentioned previously (c.f. Section 3.1), a preliminary requirement of the proposed OTT is that fiats quoted for ETH are correlated and cointegrated. If not, it indicates that trading on the bucket of Fiat/ETH pair is hardly following a mean-reversion process. We use the SPOT ETH price quoted in different USD, CAD, GBP, and EUR from 01-Jan-2018 to 01-Jan-2020 for screening and formation (Algorithm 1 line 1). Our dataset is sampled at 1min, 5min and 60min trading intervals. We want to test the stationarity of the pairs (spreads) not only for the complete formation period, but also for shorter cycles corresponding to the time frames of our arbitrage. A low correlation or cointegration result implies bad applicability of the proposed OTT. Therefore we examine the data on a rolling 1-week basis to test for correlation and stationarity via cointegration.
Table 2: Correlation and Cointegration Among Currency Pairs
Currency pair Screening Test 1m Interval 5m Interval 60m Interval
USD:CAD Correlation Cointegration 0.978 77.8% 0.998 100.0% 0.965 89.9%
USD:GBP Correlation Cointegration 0.916 66.7% 0.974 94.1% 0.908 72.5%
USD:EUR Correlation Cointegration 0.997 77.8% 0.914 72.5% 0.994 69.6%
CAD:GBP Correlation Cointegration 0.917 85.2% 0.991 72.5% 0.903 81.2%
CAD:EUR Correlation Cointegration 0.978 88.9% 0.980 94.1% 0.969 98.6%
GBP:EUR Correlation Cointegration 0.917 88.9% 0.921 98.0% 0.913 91.3%
  • This table provides average rolling correlation calculation and cointegration test results for each currency pair between 01-Jan-2018 and 01-Jan-2020.
The correlation and cointegration test results are presented in Table 2. USD:CAD under 1-week stands for Pearson’s correlation for USD:CAD between 01-Jan-2018 to 01-Jan-2020 cut into 105 weeks. The correlation we present is the simple weekly average for 105 weeks. We expect our trading target to be stationary. Therefore we also expect the spreads we trade to be cointegrated during the period we hold the positions. Similar to Pearson’s correlation test, the data are also cut into weekly slots for the Engle-Granger test. The passing percentages are presented in the table. We can see that the data generally have good correlations and cointegration relationships, so we can confidently conduct the proposed arbitrage.
5.3 Tuning Trading Parameters
Once the stationarity of the spreads is determined we compute a number of parameters including thresholds, which determine the boundaries to open and close position from the formation data (Algorithm 1 line 1).
Figure 5: Heatmap for profitability tuning under 5min-interval between 15-Jan-2020 and 01-Oct-2022
The two parameters that determine the closing and opening boundaries are Open_Threshold and Close_Threshold multiplied by the standard deviation (c.f. Table 1). Given that the assets are naturally different, and we are betting on the volatility of spreads, therefore we devise parameters based on the standard deviation. We must decide carefully on what values will be used for the trading signals. In our research, we conduct grid-search for the thresholds that are likely to generate high profitability. Figure 5 illustrates the grid search results for 5min sample interval formation dataset. It is a heatmap for the indicators of profitability (calculation method in Section 5.3.1). A darker shade means the better profit of the Open_Threshold/Close_Threshold combination. We can see that in our test phase, for 5min sampled dataset, opening at 9 standard deviations and closing at 7 standard deviations generates the best profitability. The grid search is extended to the trading intervals of 5min and 60min as well. Table 3 records the most lucrative combination for adopted trading intervals in this research.
Table 3: The most profitable trading threshold
Interval Open_Threshold Close_Threshold
1 min 11 Standard Deviations 9 Standard Deviations
5 min 9 Standard Deviations 7 Standard Deviations
60 min 7 Standard Deviations 6 Standard Deviations
  • The thresholds are calculated from retrospective grid searches from 05-Jan-2020 and 01-Oct-2022 for the most lucrative combination of Open_Threshold and Close_Threshold.
Another tunable parameter is the risk-preference parameter LAMBDA ($\lambda$), which is used to adjust the punishment factor for investment risk in Equation (2). We can tweak $\lambda$ to cater for different risk preferences. A high $\lambda$ represents a strong punishing factor for portfolio volatility, leading to a more conservative investment decision. By default, $\lambda$ is set to 1. But we will also examine the profitability of arbitrage under a riskier $\lambda$=0.5 and a more conservative $\lambda$=2.
5.3.1 Calculating P&L in base currency
Since we are trading with multiple fiat currencies, we need a unified method to measure profitability (or loss) — so-called P&L. The way we do this is to choose a base currency among participating assets — USD in our case. We obtain the foreign currency exchange rate for USD-CAD, USD-EUR, and USD-GBP from the beginning of the arbitrage testing period to the end (Table 4).
Table 4: Foreign Currency Exchange Rate for USD
Full-Cycle Market 15-Jan-2021 01-Oct-2022
USD : USD 1.000000 1.000000
USD : CAD 1.266453 1.383270
USD : GBP 0.758367 0.896861
USD : EUR 0.888224 1.019910
Bull Market 01-Jan-2021 01-Jan-2022
USD : USD 1.000000 1.000000
USD : CAD 1.273005 1.263930
USD : GBP 0.731368 0.739196
USD : EUR 0.821600 0.879569
Bear Market 01-Jan-2022 01-Jan-2023
USD : USD 1.000000 1.000000
USD : CAD 1.263930 1.354530
USD : GBP 0.739196 0.826327
USD : EUR 0.879569 0.931797
Notice that our testing dataset does not overlap with the formation period (12-Mar-2016 — 31-Dec-2021). Hence, all our arbitrage tests are on out-of-sample (never seen before) datasets.
We start with 10000 USD, CAD, GBP and EUR, each. It is equivalent to holding 42340.72 USD on 15-Jan-2021. If all the assets we hold remain the same, it will be converted as equivalent to holding 38184.03 USD on 01-Oct-2022. The assets we kept are the same, but we calculated the result as a 9.82% loss. However, it does not mean that we will naturally lose 9.82% due to no action. Since our technique tracks forex volatility as part of the trading signal, a bullish USD will navigate the arbitrage to long ETH with USD for other currencies. So we can view our profitability as a compounded effect of cryptocurrency arbitrage and forex volatility arbitrage. Another example with portfolio changes is presented in Table 5. The starting portfolio is converted to USD as stated above. The end date portfolio is converted based on end date FX. Hence we can have a comparable profitability of 42340.72 USD to 40564.62 USD as -4.19%.
Table 5: An example of profit calculation
15-Jan-2021 01-Oct-2022
USD 10000 USD = 10000 USD 12000 USD = 12000 USD
CAD 10000 CAD = 7896.06 USD 8000 CAD = 5783.40 USD
GBP 10000 GBP = 13186.23 USD 9000 GBP = 10035.00 USD
EUR 10000 EUR = 11258.42 USD 13000 EUR = 12746.22 USD
Total 42340.72 USD 40564.62 USD
5.4 Profit and Loss Results
In this section, we give the results of applying the proposed Optimal Trading Technique (OTT)(Algorithm 1 line 1) to out-of-sample dataset under different market conditions.
5.4.1 Full-Cycle market
We use the method in Section 5.3.1 to calculate the profitability in USD, presented in Table 6. The numbers in the table for various currencies are the profit (or loss) at the end of the trading period, starting with 10000 amount each. All currencies are converted to USD based on the exchange rate at the start of the trading period of 15-Jan-2021. We convert the currencies back to USD for profitability calculation on 01-Oct-2022. During this period, cryptocurrency experienced a bull market and a bear market, allowing us to test our technique through peak and trough. The results of our technique are encouraging. We achieved an annualized return of 14.87% in USD for 1min sampled dataset ($\lambda$=1). For the 5min and 60min sampled datasets, the annualized returns were 15.49% and 5.00%, which are greatly exceeding the Distance Method (DM) in all experimented cases.
Figure 6: The profitability of Full-Cycle Market between 15-Jan-2021 to 01-Oct-2022
We also conduct experiments on the conditions where investors have higher or less tolerance for portfolio volatility. We can observe that there is a clear relationship: the lower the $\lambda$, the higher the profitability. This relationship meets our expectation that the investors receive a risk premium in exchange for a more volatile portfolio.
Table 6: Annualized profitability results for $\lambda$ in full-cycle market (Adjusted)
Interval BH DM OTT($\lambda$=0.5) OTT($\lambda$=1) OTT($\lambda$=2)
1 min (TC 0.1%) 13.14% 5.33% 35.54% 14.87% 7.44%
1 min (TC 0%) 13.14% 0.73% 79.98% 40.03% 19.97%
5 min (TC 0.1%) 13.14% 0.60% 37.74% 15.49% 7.74%
5 min (TC 0%) 13.14% 4.08% 90.03% 44.98% 22.49%
60 min (TC 0.1%) 13.14% 3.93% 9.93% 5.00% 2.51%
60 min (TC 0%) 13.14% 5.54% 37.43% 18.96% 9.49%
  • 1
    The arbitrage period is 15-Jan-2021 to 01-Oct-2022. Participating assets are USD, CAD, GBP, and EUR. The absolute return is stated in percentages.
  • 2
    Abbreviations. BH=”Buy-Hold”, DM=”Distance Method”, TC=”Transaction Cost”, OTT=”Optimal Trading Technique”
  • 3
    Open_Threshold and Close_Threshold are tuned best on the most lucrative combination (Table 3)
Another comparison is the inclusion of transaction costs. We can see that even a 0.1% transaction cost significantly impacts our returns. On the surface, transaction cost charges a flat percentage of trading capital and influences the optimization process’s expectation. Our technique trades on a frequent basis; hence we can see the profitability is almost halved in three examined $\lambda$ values after including transaction costs. The Distance Method, on the other hand, does not have a significant impact due to transaction cost, because it trades less frequently (Table 10).
5.4.2 Bull market
Financial assets are subject to fluctuations, with periods of both growth and decline. However, the duration of a trading period may be one-sided. The year 2021 has been particularly noteworthy for cryptocurrencies, with Ethereum (ETH) appreciating fourfold across all major fiat currencies. Although such growth is unlikely to recur frequently, it is possible to evaluate trading techniques in extreme market conditions to determine if any specific trading traits emerge.
Figure 7: The profitability of Bull Market between 01-Jan-2021 to 01-Jan-2022
Table 7 presents the trading results for 2021. A significant growth trend is evident from the Buy-Hold strategy, which yielded a return of 412.09%. Such a circumstance is rare and occurs only in exceptional cases. Compared to full-cycle market trading, both the Distance Method (DM) and our Optimal Trading Technique (OTT)yielded more profit in the bull market. However, it was still lower than the simple Buy-Hold strategy. For a standard scenario of $\lambda=1$ and $tc=0.1\%$ over a five-minute period, the Optimal Trading Technique achieved an annual return of 36.75% (compared to 19.89% for the Distance Method under the same conditions). Single-sided markets are easier to estimate, and the expected_return and expected_risk are therefore closer to reality. The profitability trend is consistent with that of the full-cycle market. Arbitrages that do not entail transaction fees generate better returns than those with transaction fees. A lower value of Lambda ($\lambda$) implies a higher acceptance risk, resulting in a better overall arbitrage return.
Table 7: Annualized profitability results for $\lambda$ tuning in bull market
Interval BH DM OTT($\lambda$=0.5) OTT($\lambda$=1) OTT($\lambda$=2)
1 min (TC 0.1%) 412.08% 22.17% 82.13% 39.85% 19.93%
1 min (TC 0%) 412.90% 20.74% 170.04% 85.02% 42.51%
5 min (TC 0.1%) 412.08% 19.89% 77.78% 36.75% 18.38%
5 min (TC 0%) 412.90% 21.19% 165.34% 82.67% 41.34%
60 min (TC 0.1%) 412.08% 6.50% 25.23% 14.18% 6.40%
60 min (TC 0%) 412.90% 7.46% 45.25% 22.75% 11.37%
  • 1
    The arbitrage period is 01-Jan-2021 to 01-Jan-2022. Participating assets are USD, CAD, GBP, and EUR. The annualized return is stated in percentages.
  • 2
    Abbreviations. BH=”Buy-Hold”, DM=”Distance Method”, TC=”Transaction Cost”, OTT=”Optimal Trading Technique”
  • 3
    Open_Threshold and Close_Threshold are tuned best on the most lucrative combination (Table 3)
5.4.3 Bear market
We also have an interest in examining how well our technique performs in a bear market. The idea of pair trading is to pursue market neutrality, and our optimization process calculates weights to avoid including ETH in our portfolio. In theory, the process should perform as well as a bull market. Therefore, we experiment with the technique under a bear market year 2022 from 01-Jan-2022 to 31-Dec-2022. During this period, the price of an ETH dropped -61.77%. We examine the profitability with data for this price dip in Table 8.
Figure 8: The profitability of Bear Market between 01-Jan-2022 to 01-Jan-2023
During the bear market in 2022, ETH witnessed a significant drop in all major currencies. The traditional Buy–Hold strategy and most non-market-neutral techniques would experience a loss from the bear market. By not ETH at all, both Distance Method and our Optimal Trading Technique (OTT)achieved positive returns in all experimented cases. We can see from Table 8 that the patterns we found in a full-cycle market also apply here.
Table 8: Annualized profitability results for $\lambda$ tuning in bear market
Interval BH DM OTT($\lambda$=0.5) OTT($\lambda$=1) OTT($\lambda$=2)
1 min (TC 0.1%) -61.84% 1.54% 6.58% 9.57% 1.64%
1 min (TC 0%) -61.77% 1.71% 19.14% 3.29% 4.78%
5 min (TC 0.1%) -61.84% 1.86% 9.98% 4.99% 2.49%
5 min (TC 0%) -61.77% 2.08% 36.41% 18.20% 9.10%
60 min (TC 0.1%) -61.84% 5.25% 18.93% 9.01% 4.92%
60 min (TC 0%) -61.77% 5.53% 33.08% 13.93% 8.52%
  • 1
    The arbitrage period is 01-Jan-2022 to 01-Jan-2023. Participating assets are USD, CAD, GBP, and EUR. The annualized return is stated in percentages.
  • 2
    Abbreviations. BH=”Buy-Hold”, DM=”Distance Method”, TC=”Transaction Cost”, OTT=”Optimal Trading Technique”
  • 3
    Open_Threshold and Close_Threshold are tuned best on the most lucrative combination (Table 3)
5.4.4 Risk and Sharpe Ratio
Profit and risk are the two sides of the same coin. We have shown records that the lower the Lambda parameter, the higher the profit. However, we also need to avoid the possibility of capital loss. The way we measure the risk of an investment is Sharpe Ratio Sharpe (1966). Sharpe $\text{Sharpe Ratio}=S_{p}=\frac{E[R_{p}]-R_{f}}{\sigma}$, where the $R_{p}$ is the portfolio’s annualized return, $R_{f}$ is the annual risk-free rate [^3] We use 4% based on U.S. 10-Year Treasury Note at the time of writing and $\sigma$ is the annualized standard deviation of the portfolio returns.
Table 9: Position Movement for different trading strategies in markets
Full-Cycle Market Bull Market Bear Market
Buy-Hold $S_{p}$: 0.12, $\sigma$: 0.961 $S_{p}$: 1.69, $\sigma$: 1.313 $S_{p}$: -0.61, $\sigma$: 0.870
DM $S_{p}$: -0.41, $\sigma$: 0.082 $S_{p}$: 0.60, $\sigma$: 0.273 $S_{p}$: -1.51, $\sigma$: 0.019
OTT($\lambda=2$) $S_{p}$: 0.44, $\sigma$: 0.104 $S_{p}$: 2.11, $\sigma$: 0.070 $S_{p}$: -0.32, $\sigma$: 0.042
OTT($\lambda=1$) $S_{p}$: 0.68, $\sigma$: 0.193 $S_{p}$: 2.66, $\sigma$: 0.126 $S_{p}$: 0.16, $\sigma$: 0.082
OTT($\lambda=0.5$) $S_{p}$: 1.11, $\sigma$: 0.343 $S_{p}$: 4.34, $\sigma$: 0.174 $S_{p}$: 0.41, $\sigma$: 0.161
  • 1
    $S_{p}$ is Sharpe Ratio of investment portfolio calculated based on a Risk-Free annual rate of 4%, $\sigma$ is the standard deviation of investment portfolio;
  • 2
    The full-cycle market is from 15-Jan-2021 to 01-Oct-2022. The bull market is from 01-Jan-2021 to 01-Jan-2022. And the bear market is from 01-Jan-2022 to 01-Jan-2023.
  • 3
    The participating techniques are ”BH” for simple buy-and-hold; ”DM” for Distance Method; ”OCT” for our Optimal Trading technique, where $\lambda$ represents are risk control factor;
  • 4
    Portfolio USD Equivalent is a conversion of measuring the trend of multiple assets back to a unified index. We extract the exchange rate from the last trading day, and use the exchange rate to convert all assets back to USD through the full trading period;
The results from Table 9 give comparisons of different techniques. From the annualized standard deviation $\sigma$, we can see that Distance Method (DM)and Optimal Trading Technique (OTT)under all conditions have significantly lower volatility than Buy-Hold. DM is less volatile than OTT because it trades less. For OTT, there is an expected negative relation between $\lambda$ and $\sigma$. A greater value of $\lambda$ effectively reduces portfolio volatility. When measuring the overall portfolio performance with the Sharpe Ratio, lower $\lambda$ brings higher profitability, therefore, a higher Sharpe Ratio.
5.5 Strategy Indicators
To have an overview of the trading strategy, we also look into the trading metrics, not just the profitability and risk. For the consistency of measurement, one action of arbitrage is the whole process of position opening and closing. One action of trade is a one-time position opening or closing, which consists of two buy or sell orders. Therefore, the relationship can be presented in an equation: 1 Arbitrage = 2 Trades [Position Open/Close] = 4 orders [Buy/Sell].
Table 10 includes some common trading strategy indicators. It consists of three sets of experiments: Full-Cycle, Bull and Bear markets based on Section 5.4. The Full-Cycle market contains the trades from 15-Jan-2021 to 01-Oct-2022 (625 days); the bull market contains 01-Jan-2021 to 01 Jan-2022 (365 days); and the bear market contains the trades from 01-Jan-2022 to 01-Jan-2023 (365 days).
  • number of trades The total number of trades happened during the whole arbitrage period
  • % of winning The percentage of winning trades among all trades
  • % of losing The percentage of losing trades among all trades
  • % of long win The percentage of winning trades among all long trades
  • % of short win The percentage of winning trades among all short trades
  • Win/Loss ratio The ratio of winning trades against losing trades
  • Average loss The average USD loss among losing trades
  • Average Win The average USD win among winning trades
  • Largest loss The largest USD loss among losing trades
  • Largest win The largest USD win among winning trades
  • Avg holding hrs on average how many hours is an arbitrage open
We can clearly see that our technique conducts trades way more frequently than the Distance Method. It is mainly because our spread focuses on the deviation of forex movement anchored to ETH while the Distance Method concentrates on the distance. Our six experiments setup has a Win/Loss ratio of greater than 1 (Table 10), which formed the foundation of pair trading profitability. It is reasonable that the longer the interval, the less trade and, therefore, longer average holding hours. We are also glad to see that there is no particularly large value in the largest win or largest loss that could impact the overall revenue. It is a symbol of the stability of our arbitrage.
Table 10: Strategy Indicators
Indicators Fullcycle Optimal Trading Technique Distance Method
1min 5min 60min 1min 5min 60min
number of trades 13790 12166 1906 112 56 30
% of winning 53.2% 54.4% 52.6% 63.3% 53.6% 56.7%
% of losing 46.8% 45.6% 47.3% 36.7% 46.4% 43.3%
% of long win 54.8% 55.5% 52.4% 78.6% 71.4% 73.3%
% of short win 51.6% 53.2% 53.0% 48.2% 35.7% 40.0%
Win/Loss ratio 1.14 1.19 1.13 1.73 1.15 1.31
Average loss -\$6.09 -\$8.67 -\$37.31 -\$50.72 -\$174.64 -\$657.00
Average win \$7.00 \$9.21 \$37.43 \$59.04 \$210.41 \$518.33
Largest loss -\$347.31 -\$475.24 -\$1078.51 -\$299.56 -\$864.43 -\$2806.60
Largest win \$386.01 \$540.01 \$976.62 \$366.06 \$862.34 \$2690.23
Avg holding hrs 4.52 5.09 16.10 236.88 565.50 933.07
Indicators Bull Optimal Trading Technique Distance Method
1min 5min 60min 1min 5min 60min
number of trades 13490 10456 1620 400 290 122
% of winning 54.0% 55.2% 53.6% 62.5% 60.0% 58.2%
% of losing 46.0% 44.8% 46.4% 37.5% 40.0% 41.8%
% of long win 56.8% 57.1% 54.9% 74.0% 75.9% 70.5%
% of short win 51.3% 53.4% 52.3% 51.0% 44.1% 45.9%
Win/Loss ratio 1.18 1.23 1.16 1.67 1.5 1.39
Average loss -\$6.38 -\$9.61 -\$48.33 -\$53.67 -\$78.53 -\$179.77
Average win \$7.97 \$10.84 \$48.89 \$55.93 \$84.80 \$169.11
Largest loss -\$335.44 -\$804.18 -\$1502.54 -\$605.46 -\$792.19 -\$1193.86
Largest win \$359.88 \$794.02 \$1554.55 \$635.35 \$1480.52 \$1325.57
Avg holding hrs 2.88 4.03 19.83 41.99 56.79 114.67
Indicators Bear Optimal Trading Technique Distance Method
1min 5min 60min 1min 5min 60min
number of trades 2146 3306 814 48 40 32
% of winning 52.6% 52.8% 54.3% 60.4% 55.0% 53.1%
% of losing 47.4% 47.2% 45.7% 39.6% 45.0% 46.9%
% of long win 47.2% 52.8% 60.2% 50.0% 50.0% 43.8%
% of short win 58.0% 52.9% 48.4% 70.1% 60.0% 62.5%
Win/Loss ratio 1.11 1.12 1.19 1.53 1.22 1.13
Average loss -\$10.11 -\$10.40 -\$40.85 -\$85.21 -\$107.74 -\$153.77
Average win \$10.25 \$10.51 \$39.00 \$73.48 \$109.60 \$172.19
Largest loss -\$420.11 -\$737.76 -\$1342.21 -\$249.60 -\$382.87 -\$350.41
Largest win \$440.61 \$736.69 \$1387.96 \$294.02 \$435.66 \$402.41
Avg holding hrs 22.69 15.15 49.80 433.01 433.23 447.44
33footnotetext: Optimal Trading technique adopts the most lucrative Open_threshold/Close_Threshold for 1min, 5min and 60min trading interval (Table 3)33footnotetext: All the data are collected under LAMBDA as 1 and transaction cost as 0.1%
As for the comparison of the bull market and bear market, there is a significant decrease in the number of trades and a slight drop in the average profit (or loss). That is to say, although our strategy remains effective in a bear market, the quality and quantity of the arbitrage opportunities are impacted.
6 Conclusion
In this paper, we examined the feasibility and profitability of multivariate pair trading. We take fiat currencies quoted for ETH as the bucket of our pair. We construct the spread focusing on the anomaly in forex, because of the movement of forex coming from independent trades against cryptocurrencies. The trading threshold is a tunable parameter based on the standard deviation of spreads. We demonstrated a grid search result of profitability under untuned conditions. In addition to the conventional pair trading strategy, we also introduced an optimization process for the resource allocation that emerged from multivariate pair trading, which focused on the restrictions of ensuring market neutrality and utilizing most of the holding assets (without leverage).
The results are encouraging. The profitability of the proposed Optimal Trading Technique (OTT)is on average 2.72 times [^4] The average calculation in the conclusion section is calculated from 1min, 5min, 60min sampled dataset for bull, bear and full-cycle market under 0.1% transaction fee more profitable compared to the baseline approach – Distance Method. This implies the adoption of trading spread is more effective than the distance in the volatile cryptocurrency market. Under the assumption of a 4% risk-free interest rate, the Sharpe Ratio of OTT still achieved a strong 1.17. For all the tested cases, OTT achieved Win/Loss greater than 1, which is the driving force of steady capital growth.
Our bi-criteria optimization objective can be tuned with risk tolerance parameter LAMBDA ($\lambda$), which impacts the optimization process for how much risk a trader tolerates. For the full-cycle market, 5min sampled dataset, the high ($\lambda=2$), intermediate ($\lambda=1$) and low ($\lambda=0.5$) risk aversion parameter value have annualized returns of 7.74%, 15.49% and 37.74%, respectively. The corresponding Sharpe Ratio 0.44, 0.68 and 1.11 have a negative correlation with $\lambda$, which means the strong profitability compensates for the volatility brought by riskier investment choices.
Our Optimal Trading technique produces positive returns under both bull and bear market trends. In the bull market of 2021, OTT conducted 10456 trades, while in the bear market of 2022, OTT only conducted 3306 trades. Apparently the quantity and quality of arbitrage opportunities are better in a bull market.